

********************************************************************
UARS WEEKLY STATUS REPORT - 17 MAR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operation with all engineering parameters within
nominal operating range. A combination of tape recorder and TDRS coverage
has provided HALOE with near-full coverage of sunrise and sunset events.
The HALOE PI and the HALOE FOT appreciates the extra efforts by the UARS
off line engineers in developing the command loads each day to manage the
tape recorder operations to acquire the HALOE events missed by TDRS contacts.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG is working with the Flight Operations Team to
develop schedules for tape recorder operation to obtain
the maximum amount of northern hemisphere MLS data during
the next period of MLS operation.
Routine processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
7 days HAL (4.82) PROD MODE (L1-L3) (23R-18C=5L)
7 days SEN (5.0) PROD (1992) (7R-7C=0L)
22 days WIN (5.11A) PROD (L1-L3) (543R-529C-14F=0L)
5 days WIN (5.11A) LOG (5R-5C=0L)
Completed all requested days for WINDII V5.11A mission
reprocessing.
Continue to process a group of days that are needed by HALOE
in support of the SOLVE Mission. (UARS DAYS 3083 - 3105)
The CDHF will load January 2000 current day DAAC data in March 2000.
UARS CDHF REHOST:
- Successful processing on Rehost System.
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM.
- Testing HRDI V4.4, SOLSTICE V4.62, and WINDII V5.11A.
***********
FOT
***********
REPORTING PERIOD
This report covers 04 March, 2000 (Orbit 46338, GMT Day 064) through 10
March 2000 (Orbit 46442, GMT Day 070).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 064/0829 46343 Off
HALOE 064/1014 46344 On
PEM 065/0036 46353 AXIS-2 Off
WINDII 065/0036 46353 On
HALOE 065/2115 46366 Off
HALOE 065/2220 46367 On
The beta angle ranged between the angles of 50.3 degrees to 41.5 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+2.0 degrees on 21 March 2000 (DOY 081).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery
1 remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and
Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data for
each battery on 03 March 2000 and 07 March 2000 are provided in Appendix D.
The clock error ranged from -11.4 to 10.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT = 22
% HALOE = 04
% PEM = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/021 - 2287.496849
- Transponder B: 00/060 - 2287.499207
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/062 show the observatory data loss to be 565 hours,
44 minutes, 32.601 seconds (an increase of 0 minutes 0 seconds since the last
report period).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT spent most of the week working on the UARS OBC safehold Failure
Detection and Correction algorithms. Three TMONs have been identified for
complete OBC attitude anomaly protection. The algorithms have been completed.
In addition, the Relative Time Sequence (RTS) has been written which will
command UARS to Earth Pointing safehold in an OBC controlled anomaly. The
remaining work is as follows
- Code TMONs and compile using OBCSTF utility.
- Code RTSs for use in the UARS Simulator.
- Meet with simulator support and update test scenarios.
- Test TMONs and RTSs for coding errors and system response.
- Submit for Peer Review.
The FOT finally located the algorithms for the elusive telemetry pseudos.
Global read-only privileges were temporarily granted to the FOT so that a
global search could be done. These pseudos were needed for some of the
safehold work being done as well as for the IMOC task.
MEETINGS
No meetings this week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for March 21, 2000 (DOY 081). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 March 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46338 064/0032 00-083 OBC OBC EXEC CLEAR
46340 064/0219 00-084 OBC SOFTWARE DUMP
46342 064/0549 00-081 NBTR NBTR OPS
46343 064/0829 00-085 HALOE POWER OFF HI TEMP
46344 064/1014 00-086 HALOE POWER ON
46349 064/1751 00-081 NBTR NBTR OPS
46349 064/1751 00-087 NBTR NBTR OPS
46353 065/0036 00-088 PEM AXIS - 2 POWER OFF
46353 065/0036 00-089 WINDII POWER ON
46357 065/0654 00-087 NBTR NBTR OPS
46358 065/0732 00-090 OBC Clk Adj fr 20.6 to 0.00 msec/day
46358 065/0832 00-087 NBTR NBTR OPS
46365 065/1933 00-091 NBTR NBTR OPS
46366 065/2115 00-092 HALOE POWER OFF
46367 065/2220 00-093 HALOE POWER ON
46373 066/0734 00-091 NBTR NBTR OPS
46373 066/0835 00-091 NBTR NBTR OPS
46380 066/1854 00-095 NBTR NBTR OPS
46387 067/0557 00-094 OBC Clk Adj fro 0.00 to 20.6 msec/day
46388 067/0737 00-095 NBTR RTS - 58
46388 067/0838 00-095 NBTR NBTR OPS
46394 067/1758 00-096 NBTR NBTR OPS
46409 068/1800 00-097 NBTR NBTR OPS
46416 069/0519 00-098 OBC Clk Adj fr 20.6 to 0.00 msec/day
46417 069/0701 00-097 NBTR NBTR NOOPS
46418 069/0846 00-097 NBTR RTS - 58
46435 070/1110 00-099 NBTR NBTR OPS
46439 070/1805 00-100 NBTR NBTR OPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 March 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 46436
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None to report this period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 March 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46358 065/07:35 (05 Mar 00) from 20.6 to 0.00 msec/day
46387 067/06:26 (07 Mar 00) from 0.00 to 20.6 msec/day
46416 069/05:34 (09 Mar 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 March 2000
03 Mar 2000 - GMT Day 063 Beta = 50.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ------ ------- -----
1 -56.0/-78.4 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +33.6/-11.2 +5.47/+3.87 +26.4/-15.6 25.4 1.050 20.1
3 +33.6/-11.2 +4.51/+2.60 +25.6/-15.2 25.3 1.048 19.9
07 Mar 2000 - GMT Day 067 Beta = 47.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------ ------- ------
1 -50.4/-72.8 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +22.4/-22.4 +5.47/+4.51 +27.2/-15.2 24.8 1.042 21.3
3 +44.8/-5.6 +4.51/+3.23 +26.8/-15.6 24.8 1.040 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 March 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-005 FOT 46361 065/1237 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No closure information received
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-005 FOT see list March 00 14170 Generic Late Acq.
SEVNTFW00/10